OpenStudy (anonymous):

ONLY HAVE 2 QUESTIONS XDD

OpenStudy (anonymous):

Which set of side lengths defines a right triangle? A. 4, 8, 9 B. 5, 9, 10 C. 6, 10, 13 D. 7, 24, 25

OpenStudy (anonymous):

Just use the Pythagorean theorem

OpenStudy (anonymous):

i came here for help yo XD

OpenStudy (anonymous):

i already did i got C

OpenStudy (anonymous):

And I'm helping you. C would be wrong. How did you get C?

OpenStudy (anonymous):

I MEAN NOT C. B sorry.

OpenStudy (anonymous):

How did you get B then?

OpenStudy (anonymous):

a^2 + b^2 = C^2

OpenStudy (anonymous):

Now plug in the numbers in B for a b and c

OpenStudy (anonymous):

i still think its B XD

OpenStudy (anonymous):

So you think 25+81=100?

OpenStudy (anonymous):

oh....A

OpenStudy (anonymous):

And why do you say A?

OpenStudy (anonymous):

bc im guessing at this point XD

OpenStudy (anonymous):

Well plug in the values in A for a b and c and tell me what you get.

OpenStudy (anonymous):

its either A or D

OpenStudy (anonymous):

So.. XD

OpenStudy (anonymous):

Plug in the values in A for the Pythagorean theorem.

OpenStudy (anonymous):

finee.

OpenStudy (anonymous):

Yeah its A

OpenStudy (anonymous):

Did you plug in the values? What did you get?

OpenStudy (anonymous):

wait lemme try again o.o

OpenStudy (anonymous):

4^2+8^2+9^2?

OpenStudy (anonymous):

No, it would be \[4^2+8^2=9^2\]

OpenStudy (anonymous):

161

OpenStudy (anonymous):

Use the formula I just gave you. You do not add all three together.

OpenStudy (anonymous):

lmao

OpenStudy (anonymous):

80 = 81 WHAT XD

OpenStudy (anonymous):

Yeah, it doesn't so that means A is wrong

OpenStudy (anonymous):

okay D thanks one more XD

OpenStudy (anonymous):

What is the value of x? Round to the nearest tenth. cm http://static.k12.com/calms_media/media/1582000_1582500/1582242/1/d312962d0705dffa9b47eba97a44e3dcfceee918/MS_IMC-150209-150419.jpg

OpenStudy (anonymous):

you would use Pythagorean theorem again

OpenStudy (anonymous):

10^2 + 12^2 = x^2

OpenStudy (anonymous):

224 =

OpenStudy (anonymous):

and then you would square root 224

OpenStudy (anonymous):

15.6

OpenStudy (anonymous):

no its about 14.9

OpenStudy (anonymous):

noo o.o

OpenStudy (anonymous):

google that ishhh

OpenStudy (anonymous):

i did lol

OpenStudy (anonymous):

its 14.966

OpenStudy (anonymous):

ur right

OpenStudy (anonymous):

XD

OpenStudy (anonymous):

lol i know :P